evening 4 Oct 99

Marj Corcoran (corcoran@physics.rice.edu)
Mon, 4 Oct 1999 23:49:45 -0500 (CDT)

Shift Summary for Evening shift, 4 Oct 99

On Shift: M. Corcoran and Leo Bellantoni

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs notv included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Beam off all shift due to CHL problems


Runs ended this shift:
Run 14802--cosmics, not to tape


Tapes brought to FCC this shift: None


Description of shift: No beam.


run spill time comment
14802 13:30 Evening shift arrives. Problems with CHL continue.
Hogan says "don't touch anything". Shutter is still up,
no one has even made a controlled access.

Hogan discovered that the new Xt's and T0's which
he thought had been installed over the weekend in fact
had not been installed. They are now installed,really.

Arun and Marj look over L3 code for trigger 7

Leo and Sasha discuss the TRD's

Marj, Leo, and RickT discuss mm0kine. Leo promises
to change the name in the pi0 mu mu paper.

19:30 Marj and Leo, in desperation, look over the Mojave
Desert phone booth web page. Some people have way too
much time on their hands including, apparently, us.

20:20 Gas check list done. Mail sent to DNorth to remind him
that an alcohol fill is due.

22:25 KTEVANA errors; KFREND. Still taking cosmics

23:00 Marj creates kumac which looks at DC1 current per
spill, plots the distribution, and does a fit. This
plot quantifies the "few hertz" spill quality
problem. See Log XXII, pg 59-60.

23:30 Jason L arrives early for his beamless shift. Does not
bring knitting, although he had been advised to do so.


Items to follow up on next shift: None that we know of.